Estimate the Hall field and magnetic field for following case: dimension of a rectangular sample of
semiconductor: width – 2mm, thickness – 1mm, Hall coefficient for the sample - 10-2 m3
/C. On passage of
1 mA current through the sample, 1mV Hall voltage is developed.
1
Expert's answer
2020-08-31T12:47:26-0400

The output Hall voltage is


"V_H=R_H\cdot\frac{IB}{t},\\\space\\
B=\frac{V_Ht}{R_HI}=\frac{0.001\cdot0.001}{10^{-2}\cdot0.001}=0.1\text{ T}."

Hall field is


"E_H=\frac{V_H}{w}=\frac{0.001}{0.002}=0.5\text{ V/m}."